### v3

@@ 3.2 Methodology
Formally, our data set comprises $n=161$ characters ($i=1,\dots,n$) observed over $T=194$ days ($t=1,\dots,T$). $T^{pre}$ is the length of the period (150 days) before the coming-out event, which occurs at time $T^{pre}+1$ (i.e., June $1^{st}$, $2022$). For each unit $i$ and time $t$, we denote the observed pick rate as $Y_{i,t}$. We represent the coming out as a binary variable $C_{i}\in\left\{0,1\right\}$ equal to one if character $i$ discloses his sexual orientation at time $T^{pre}+1$. We then posit the existence of two potential pick rates $Y_{i,t}^{c}$, where one denotes the pick rate in the absence of disclosure ($Y_{i,t}^{0}$) and the other denotes the pick rate in the presence of disclosure ($Y_{i,t}^{1}$).242424 These potential outcomes are based on Rubin’s model for causal inference [139].
@@
Since Graves’ sexual orientation has been disclosed after period $T^{pre}$, under a standard SUTVA assumption [[, see, e.g.,]]imbens2015causal we observe $Y_{1,t}=Y_{1,t}^{1}$ for all $t>T^{pre}$. Thus, as shown in equation (1), the challenge in estimating our causal effects of interest is to estimate $Y_{1,t}^{0}$ for $t>T^{pre}$, i.e., how Graves’ pick rates would have evolved in the absence of the disclosure. To this end, we can construct a synthetic control unit that approximates the pick rates of Graves before the coming out. The idea is that if the synthetic control and Graves behave similarly before the disclosure, then the synthetic control can serve as a valid counterfactual.
@@
$\begin{gathered}\hat{\omega}=\operatorname*{arg\,min}_{\omega\in\Omega}\ell\left(\omega\right),\\[4.30554pt] \ell\left(\omega\right)=\sum_{t=1}^{T^{pre}}\left(\sum_{i=2}^{n}\omega_{i}Y_{i,t}-Y_{1,t}\right)^{2}+\zeta^{2}T^{pre}\|\omega\|_{2}^{2},\quad\Omega=\left\{\omega\in\operatorname{\mathbb{R}}_{+}^{n-1}:\sum_{i=2}^{n}\omega_{i}=1\right\},\end{gathered}$ (2)
@@
We employ the “placebo approach” of [179] to estimate the variance of $\hat{\tau}$. We then use the estimated variance to construct asymptotically valid conventional confidence intervals.252525 The validity of this placebo approach hinges on a homoskedasticity assumption which requires that treated and control units have the same noise distribution. In general, with only one treated unit, nonparametric variance estimation for treatment effect estimators is typically impossible without a homoskedasticity assumption [179].
4 Results
In this section, we present our main results. First, we present our main findings. We then investigate heterogeneity across players with different levels of in-game performance. Finally, we explore the possibility of regional variations in attitudes toward the LGB community by replicating our analysis across different macro-regions. Appendix B provides a series of robustness checks that support the reliability of our estimates.
4.1 Main Results
We apply the synthetic control estimator of Section 3.2 to estimate the effects of the coming-out event on players’ revealed preferences for Graves. To mitigate the potential for spillover effects, we exclude Twisted Fate and other four characters (Diana, Leona, Nami, and Neeko) that were already members of the LGB community prior to the coming-out event from the donor pool.262626 Nevertheless, even if included in the donor pool, the estimator assigns these characters zero weight.
@@ 4.1 Main Results
Figure 2: Graves’ daily pick rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 2 displays Graves’ actual and synthetic pick rate series.272727 Figure A.II in Appendix A displays the identities and the contributions of the characters in the donor pool with non-zero estimated weights. Overall, our analysis suggests a substantial negative impact of the coming-out event on players’ preferences for Graves. Before the disclosure, the synthetic control estimator closely approximates the trajectory of Graves’ pick rates, providing support for the estimator’s ability to predict the counterfactual series. However, starting from June $1^{st}$, $2022$, the two series diverge substantially, with Graves’ pick rates consistently dropping below those of the synthetic control. This gap persists over time, extending even beyond the conclusion of LGBT Pride Month. The estimated average effect, reported in the first column of Table B.I in Appendix B, is 7.60-7.60 percentage points and is statistically different from zero at the $5\%$ significance level.282828 To provide further evidence that our main estimate is not driven by chance, we reassign the treatment to all control characters in our data set. This allows us to generate a distribution of “placebo” effects, which serves as a basis for comparing the impact on Graves [[, see, e.g.,]]abadie2010synthetic, abadie2015comparative. We report the ratios of post- to pre-treatment root mean squared error (RMSE) in Figure A.III in Appendix A, where Graves ranks third in terms of RMSE ratio. This implies that the impact on Graves is unusually large compared to the distribution of placebo effects, reinforcing the interpretation that our analysis provides significant evidence of a negative effect of the coming-out event. This implies a decline of $40.98\%$ of the pre-treatment average preferences for Graves.
We conduct a series of robustness checks, detailed in Appendix B, all of which reinforce our main finding. First, our results are robust to alternative estimation strategies and donor pool compositions: we obtain consistent estimates using a regularized synthetic control estimator and when restricting the donor pool to characters unlikely to act as substitutes for Graves. Second, we assess predictive accuracy by artificially backdating the treatment [[, see, e.g.,]]abadie2022synthetic; discrepancies between the actual and synthetic series emerge only on the true disclosure date and closely match those in the main analysis, demonstrating the estimator’s ability to forecast the counterfactual series and supporting a no-anticipation assumption. Third, a leave-one-out analysis [[, see, e.g.,]]abadie2021using confirms that no single donor character drives the results.
4.2 Player Heterogeneity
To examine whether responses to the coming-out event differ by player ability, we conduct a heterogeneity analysis based on a proxy for player skill derived from pre-treatment in-game performance. Specifically, we compute each player’s win rate prior to the treatment date, restricting the sample to the same 7,4217{,}421 users selected for our player-level data set to ensure the reliability of this metric. We then classify these players into above- and below-median skill groups based on this metric. Using this stratification, we replicate our main synthetic control analysis separately for each group. The estimation proceeds identically to our baseline specification of Section 4.1.
This stratification allows us to investigate whether the decline in Graves’ pick rate is concentrated among particular subpopulations—such as lower-skilled or more casual players—or whether it reflects a broader behavioral response. Consistent estimates across skill groups would strengthen the external validity of our findings, mitigating concerns that the observed response is driven by a narrow or non-representative segment of the player base.
Figure A.IV in Appendix A plots the distribution of overall win rates and total matches for players in each skill group. Although the win rate distributions overlap substantially, they differ in their means, with players classified as higher-skilled exhibiting slightly higher win rates on average. In contrast, the distribution of total matches is nearly identical across groups, indicating that conditioning on pre-treatment win rate does not mechanically segment players by engagement level. This is important for interpretation: it implies that our stratification holds constant (approximately) overall game activity while isolating variation in a meaningful proxy for player quality.
(a) Players below median skill
(b) Players above median skill
Figure 3: Graves’ daily pick rates and synthetic control estimation results by skill group. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 3 presents the results. We observe a sustained decline in Graves’ pick rate immediately following the coming-out event in both the below- and above-median skill groups. The timing and persistence of the divergence from the synthetic counterfactual are remarkably similar across the two groups and closely mirror the patterns documented in our pooled specification.292929 Appendix B shows that these findings are robust to backdating and leave-one-out exercises. These findings indicate that the estimated effect is not concentrated among players of a particular ability level but instead reflects a widespread behavioral response, thereby enhancing the external validity of our findings and mitigating concerns that the observed decline in Graves’ popularity is driven by a narrow or non-representative subset of users.303030 While qualitatively similar, the pre-treatment pick rates of Graves are lower in both skill groups relative to the main specification. This difference arises from the additional sample restriction imposed here, whereby we include only players who participated in at least 10 matches both before and after the coming-out event to ensure a meaningful proxy for skill. The main specification, by contrast, includes the full player base, potentially encompassing more casual users with higher Graves usage. As such, differences in baseline levels should be interpreted as compositional.
4.3 Regional Heterogeneity
Previous research has demonstrated that attitudes toward LGB people can substantially vary between countries, causing also a large variation in the number of individuals openly identifying with the LGB community [[, see, e.g.,]]badgett2020economic, badgett2021lgbtq.313131 An OECD report shows that even within a set of relatively comparable OECD countries, the size of the LBG communities differs by a factor of four [171]. To explore regional differences in players’ attitudes towards this community, we exploit the geographic information associated with each match. Specifically, we divide the sample according to the macro-regions defined in Section 3.1 and apply the synthetic control estimator described in Section 3.2 separately within each region.
Figure 4: Graves’ daily pick rates and synthetic control estimation results by region. The actual series are smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Figure 4 displays the results.323232 Figure A.V in Appendix A displays the identities and the contributions of the characters in the donor pool with non-zero estimated weights. The synthetic control estimator closely approximates the trajectory of Graves’ pick rates for matches in Europe and Latin America before the disclosure, exhibiting pre-treatment root mean squared errors comparable to those of the pooled specifications. However, discrepancies arise in Korean and North American matches, where the pre-treatment root mean squared error is two to three times higher than that achieved with European and Latin American matches. In Europe, Korea, and Latin America, we estimate negative and persistent effects of the coming-out event on players’ preferences for Graves. Point estimates for the average treatment effect in these regions, reported in Table A.II in Appendix A, are consistently negative and are statistically significant at the 5% level across most of the considered specifications. In North America, where the pre-treatment root mean squared is relatively large, we are unable to determine the sign of the impact, as the confidence intervals for the estimated average effect consistently encompass zero.
The estimated average effect varies substantially across regions, with the largest effect observed in Europe ($62.23\%$ relative to the pre-treatment average preferences) and the smallest effect observed in Korea ($31.82\%$) in our main specification, indicating heterogeneous responses to the coming-out event. However, it is important to consider that the different magnitudes of the estimated effects do not necessarily reflect differential attitudes towards the LBG community, as various other factors may differ across servers. One such factor could be the differential levels of competitiveness on different servers, which may affect the character selection process by introducing different levels of subjectivity. We therefore carefully interpret our results in this section as evidence for anti-LGB sentiments in various regions, rather than interpreting the differences in the effect sizes.
5 Mechanisms
@@
In Section 4, we established evidence of a substantial negative impact of the coming-out event on players’ revealed preferences for Graves. However, the players’ decision to switch from this character might be influenced by factors beyond preferences for LGB status. The objective of this section is to eliminate these alternative channels, thereby enhancing the plausibility of social stigma as the primary explanation for the observed behavior.
First, we examine the idea that shifts in character relative strengths could explain our estimated effect. We rule out this possibility in Section 5.1 by demonstrating that Graves’ strength remained unaffected by the coming-out event. Second, we explore the potential influence of players’ skills on their decision to abandon Graves. In Section 5.2, we show that players’ skills have no correlation with the choice to drop the character, thus dismissing the possibility that gameplay factors are the driving force behind the players’ observed behavior. Third, we investigate whether players transitioning away from Graves experience any performance-related consequences. This is the topic of Section 5.3, where we present evidence that switching to other characters does not affect the performance of the players involved. This emphasizes our ability to measure players’ true social attitudes and stigma, avoiding any potential biases stemming from strategic performance considerations. Fourth, in Section 5.4, we dismiss the possibility that the release of a new potentially substitute character drives the players’ decisions to switch away from Graves.
Finally, we acknowledge that questions may arise about whether the findings of Section 4 are solely a consequence of Graves’ disclosure or if they are influenced by the broader context of LGBT Pride Month. We exploit the presence of other playable characters with sexual minority status to show that LGBT Pride Month is unlikely to explain our findings. Specifically, Appendix D introduces a theoretical framework that formalizes the existence of two “simultaneous treatments” and outlines sufficient assumptions that enable us to separate the impacts of coming out and LGBT Pride Month on players’ preferences for Graves. The results, detailed in Section 5.5, support the interpretation that the estimated effects are driven by Graves’ disclosure.
5.1 Graves’ Strength
@@
Crucial to the plausibility of social stigma attached to playing an LGB character as the primary explanation for the players’ observed behavior is the fact that Graves’ strength remained unaffected by the coming-out event, as any change in character relative strengths could explain why players’ preferences shift away from Graves.
@@
To address this concern, we employ the synthetic control estimator described in Section 3.2 to examine the potential impact of the coming-out event on Graves’ strength. We measure characters’ strength using daily win rates.
@@
Figure 5: Graves’ daily win rates and synthetic control estimation results. The actual series is smoothed by a Nadaraya-Watson regression before applying the synthetic control estimator. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
@@
Figure 5 displays the results. Overall, our analysis reveals that the coming-out event had no impact on Graves’ strength. Despite the actual series exhibiting daily fluctuations around the $50\%$ mark, the synthetic control estimator effectively captures its pre-treatment trend, showcasing its ability to predict the counterfactual trend. After the treatment date, the synthetic control estimator continues to align with Graves’ win rate trend, confirming that the character’s strength was unaffected by the disclosure. The average effect is estimated to be 1.046-1.046 percentage points (standard error: $5.298$), and the conventional $95\%$ confidence interval encompasses zero, indicating a failure to reject the null hypothesis of no effect. These findings demonstrate that Graves’ strength remained unchanged during the coming-out event, dismissing the possibility of a shift in his strength as an explanation for the results of Section 4.
@@
Moreover, we note that players have real-time access to detailed information regarding characters’ strengths, weaknesses, and overall performance, as numerous websites continuously provide updated data on characters’ in-game statistics.333333 Examples of such websites include https://lolalytics.com/lol/graves/build/ and https://www.leagueofgraphs.com/champions/stats/graves. Therefore, players were well-informed that no game-relevant skills or attributes were altered during the treatment period, and they could observe that Graves’s strength remained consistent. These factors suggest that the negative impact of the coming-out event estimated in Section 4 is unlikely to be driven by actual or presumed changes in character relative strengths.
5.2 Players’ Skills
If highly skilled players exhibit distinct preferences for Graves or are less influenced by the character’s sexual orientation, the decision to switch from Graves might be driven by gameplay factors rather than social preferences for sexual orientation, thus challenging our social stigma narrative.
@@ 5.2 Players’ Skills
To address this concern, we examine the correlation between players’ skills and their decision to abandon Graves. We classify players into two groups based on their preferences for Graves before his disclosure: the first group comprises those who chose Graves in at least $5\%$ of their matches before the coming-out event (henceforth labeled as prior users, $n=953$), while the second group comprises the remaining players (henceforth labeled as non-prior users, $n=6468$). We then examine performance differences both within and between these groups before and after the treatment.
@@
Figure 6: Players’ average pick rates for Graves and performance measures. Players are divided into two groups based on their preferences for Graves before his disclosure. The panels display the average pick rates for Graves, number of daily matches, and win rate of each group before and after the coming-out event.
@@
The top panel of Figure 6 displays the average pick rate for Graves among prior and non-prior users before and after the treatment. We observe a sharp decline in pick rates among prior users following the coming-out event, similar in size to the decrease shown in Figure 2 and Table 1. Conversely, non-prior users exhibit a marginal increase in average pick rates post-treatment, although this increase is practically negligible.
@@
In the remaining panels of Figure 6, we investigate whether prior and non-prior users differ in their in-game behavior. First, the bottom left panel shows the average number of daily matches played by each group. We observe similar engagement levels across groups, both before and after the treatment, indicating that prior and non-prior users tend to play a comparable number of matches per day. Notably, neither group appears to disengage from the game following Graves’ disclosure. On the contrary, the number of daily matches slightly increases post-treatment for both groups, suggesting that players are not leaving the game but instead redirecting their attention to other characters. Supporting this interpretation, Figure A.VI in Appendix A shows no meaningful change in role preferences for either group following the event. This pattern is consistent with the idea that players face low switching costs and have access to suitable alternatives, enabling them to express their preferences without disrupting their gameplay. These results align with those displayed in Table 1.
Second, the bottom right panel displays the players’ average win rates. We observe no substantial disparities within and between groups, indicating that the preference for Graves and the decision to abandon this character are unrelated to players’ skill levels. Overall, these findings dismiss the possibility that game-play factors are the driving force behind the estimated effects of Section 4, lending additional support to the social stigma attached to playing an LGB character as the mechanism underlying the players’ observed behavior.
5.3 Players’ Performance

### v4

@@ 6 Conclusion
Formally, our data set comprises $n=161$ characters ($i=1,\dots,n$) observed over $T=194$ days ($t=1,\dots,T$). $T^{pre}$ is the length of the period (150 days) before the coming-out event, which occurs at time $T^{pre}+1$ (i.e., June $1^{st}$, $2022$). For each unit $i$ and time $t$, we denote the observed pick rate as $Y_{i,t}$. We represent the coming out as a binary variable $C_{i}\in\left\{0,1\right\}$ equal to one if character $i$ discloses his sexual orientation at time $T^{pre}+1$. We then posit the existence of two potential pick rates $Y_{i,t}^{c}$, where one denotes the pick rate in the absence of disclosure ($Y_{i,t}^{0}$) and the other denotes the pick rate in the presence of disclosure ($Y_{i,t}^{1}$).373737 These potential outcomes are based on Rubin’s model for causal inference [48].
@@
Since Graves’ sexual orientation has been disclosed after period $T^{pre}$, under a standard SUTVA assumption [[, see, e.g.,]]imbens2015causal we observe $Y_{1,t}=Y_{1,t}^{1}$ for all $t>T^{pre}$. Thus, as shown in equation (B.1), the challenge in estimating our causal effects of interest is to estimate $Y_{1,t}^{0}$ for $t>T^{pre}$, i.e., how Graves’ pick rates would have evolved in the absence of the disclosure. To this end, we can construct a synthetic control unit that approximates the pick rates of Graves before the coming out. The idea is that if the synthetic control and Graves behave similarly before the disclosure, then the synthetic control can serve as a valid counterfactual.
@@
$\begin{gathered}\hat{\omega}=\operatorname*{arg\,min}_{\omega\in\Omega}\ell\left(\omega\right),\\[4.30554pt] \ell\left(\omega\right)=\sum_{t=1}^{T^{pre}}\left(\sum_{i=2}^{n}\omega_{i}Y_{i,t}-Y_{1,t}\right)^{2}+\zeta^{2}T^{pre}\|\omega\|_{2}^{2},\quad\Omega=\left\{\omega\in\operatorname{\mathbb{R}}_{+}^{n-1}:\sum_{i=2}^{n}\omega_{i}=1\right\},\end{gathered}$ (B.2)
@@
We employ the “placebo approach” of [5] to estimate the variance of $\hat{\tau}$. We then use the estimated variance to construct asymptotically valid conventional confidence intervals.383838 The validity of this placebo approach hinges on a homoskedasticity assumption which requires that treated and control units have the same noise distribution. In general, with only one treated unit, nonparametric variance estimation for treatment effect estimators is typically impossible without a homoskedasticity assumption [5].
Supplementary Information C Robustness Checks
We examine the robustness of our main results from Section 4.1 to alternative estimation strategies and variations in the composition of the donor pool. In particular, we repeat our analysis employing a regularized synthetic control estimator (see Section B) and explore different donor pool configurations focusing on characters from distinct roles. Notably, Graves is predominantly designed for and played in three of the possible roles within a team. Consequently, there is a possibility of spillover effects on other characters mainly played in these positions, as players transitioning away from Graves are likely to switch to these alternatives.393939 Table 1 and the findings of Section 5.2 support this intuition. To mitigate this potential for spillover effects, we restrict our donor pool to characters that are “non-substitutes” of Graves, that is, those primarily designed for the remaining two roles.404040 Graves is predominantly designed for and played in the top, jungle, and mid positions. Therefore, we consider characters designed for and played in the bottom and support positions as “non-substitutes.”
Table C.I: Main results.
Synthetic Controls Regularized Synthetic Controls (1) (2) (3) (4) All characters Only non-substitutes All characters Only non-substitutes $\hat{\tau} -$7.112 -6.922 -5.462 -5.28 $p\text{-value}$ 0.005 0.071 0.027 0.17 $95\%$ CI [-12.094, -2.130] [-14.425, 0.580] [-10.289, -0.636] [-12.837, 2.269] N. Donors 4 4 6 6 RMSE 2.734 2.875 2.770 2.877 Pre-treatment average 18.557 18.557 18.557 18.557
Notes. Point estimates, $p$-values, and $95\%$ confidence intervals for $\hat{\tau}$. Additionally, the number of donors receiving a non-zero weight and the pre-treatment root mean squared error are displayed. Each column corresponds to a different specification, with the specifications differing solely in the employed estimator and donor pool composition.
Table C.I displays the results. For any donor pool composition, the results are not sensitive to the choice of the regularization parameter, and point estimates are consistently negative. The results are consistent also quantitatively across all specifications, with a decline ranging between $40.98\%$ and $38.17\%$ of the pre-treatment average preferences for Graves. Overall, these results support our main finding of a substantial negative impact of the coming-out event on players’ preferences for Graves.
Figure C.I: Robustness checks for main results. The upper panel shifts the coming-out event ten days earlier, with the new treatment date denoted by the vertical gray dashed line. The lower panel reports leave-one-out estimates of the synthetic control series, obtained by excluding one of the characters of Figure A.III at a time from the donor pool.
We also assess the credibility of the synthetic control estimator by conducting a robustness check that artificially shifts the coming-out event ten days earlier. This backdating exercise allows us to evaluate the estimator’s predictive accuracy during a ten-day hold-out period [[, see e.g.,]]abadie2022synthetic. The upper panel of Figure C.I presents the results of this analysis. We observe three key findings. First, the estimated effects remain qualitatively and quantitatively consistent, confirming a negative and persistent impact of the coming-out event on players’ revealed preferences for Graves. Second, the synthetic control estimator demonstrates a good fit during the hold-out period, indicating its ability to accurately capture Graves’ usage prior to the disclosure. Third, the actual and the synthetic series begin to diverge on the true day of disclosure, even when the estimator has no knowledge of the actual disclosure date. The absence of estimated effects before the coming-out event also lends support to the plausibility of a no-anticipation assumption [[, see e.g.,]]abadie2021using.
We conduct an additional robustness test by performing a leave-one-out exercise, where we repeatedly estimate the synthetic control series by excluding one character with non-zero estimated weights at a time from the donor pool [[, see e.g.,]]abadie2021using. The lower panel of Figure C.I presents the results of this analysis. Overall, our finding of a negative and persistent impact of the coming-out event on players’ preferences for Graves is robust to the exclusion of any particular character. Most of the leave-one-out synthetic series closely align with the main estimate, thus reinforcing the robustness of the main conclusion of our study. One leave-one-out series falls beneath the other synthetic series, suggesting a somewhat reduced, although still negative, impact. However, this series diverges from the actual series in the weeks prior to the treatment, which undermines the reliability of its results.414141 This series is obtained by excluding the character Ezreal from the donor pool. Figure A.III in Supplementary Information A shows that this character receives the largest weight in our main specification. Therefore, the divergence of this leave-one-out series from the actual series is unsurprising.
Finally, we implement a within-role placebo design by replicating the synthetic control analysis for the four most frequently selected jungle characters (excluding Graves) in the pre-treatment period. These characters share the same primary role as Graves and are therefore exposed to similar role-specific dynamics and potential demand shifts. The results are reported in Figure C.II. None of these placebo units exhibits a sharp and sustained post-disclosure divergence comparable to that observed for Graves. In particular, we do not observe a synchronized decline in jungle pick rates around June 1, 2022. While some champions display idiosyncratic fluctuations, these movements are not aligned with the disclosure date and are not persistent. Overall, these patterns indicate that the estimated effect for Graves is not driven by broader shocks to jungle demand or role-specific changes, but is instead character-specific.
@@
Figure C.II: Top four (excluding Graves) jungle characters’ daily pick rates and synthetic control estimation results. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
Supplementary Information D Mechanisms
@@
In Section 4, we established evidence of a substantial negative impact of the coming-out event on players’ revealed preferences for Graves. However, the players’ decision to switch from this character may reflect factors other than a response to the disclosure itself. The objective of this section is to eliminate these alternative channels, thereby strengthening the interpretation that the observed behavior reflects a response to the disclosure rather than confounding gameplay or design factors.
D.1 Graves’ Strength
@@
Crucial for interpreting the estimated effects as responses to the disclosure is the fact that Graves’ strength remained unaffected by the coming-out event, as any change in character relative strengths could explain why players’ preferences shift away from Graves.
@@
To address this concern, we employ the synthetic control estimator described in Section B to examine the potential impact of the coming-out event on Graves’ strength. We measure characters’ strength using daily win rates.
@@
Figure D.I: Graves’ daily win rates and synthetic control estimation results. The dashed vertical line denotes the day of disclosure, and the rainbow area highlights LGBT Pride Month.
@@
Figure D.I displays the results. Overall, our analysis reveals that the coming-out event had no impact on Graves’ strength. Despite the actual series exhibiting daily fluctuations around the $50\%$ mark, the synthetic control estimator effectively captures its pre-treatment trend, showcasing its ability to predict the counterfactual trend. After the treatment date, the synthetic control estimator continues to align with Graves’ win rate trend, confirming that the character’s strength was unaffected by the disclosure. The average effect is estimated to be 1.311-1.311 percentage points (standard error: $4.545$), and the conventional $95\%$ confidence interval encompasses zero, indicating a failure to reject the null hypothesis of no effect. These findings demonstrate that Graves’ strength remained unchanged during the coming-out event, dismissing the possibility of a shift in his strength as an explanation for the results of Section 4.
@@
Moreover, we note that players have real-time access to detailed information regarding characters’ strengths, weaknesses, and overall performance, as numerous websites continuously provide updated data on characters’ in-game statistics.424242 Examples of such websites include https://lolalytics.com/lol/graves/build/ and https://www.leagueofgraphs.com/champions/stats/graves. Therefore, players were well-informed that no game-relevant skills or attributes were altered during the treatment period, and they could observe that Graves’s strength remained consistent. These factors suggest that the negative impact of the coming-out event estimated in Section 4 is unlikely to be driven by actual or presumed changes in character relative strengths.
D.2 Engagement and Performance Across Usage Groups
If the observed decline in Graves’ usage were driven by gameplay considerations rather than the disclosure itself, we would expect systematic differences in engagement or performance across players with different pre-disclosure usage patterns.
@@ D.2 Engagement and Performance Across Usage Groups
To address this concern, we compare engagement and performance outcomes across players with different pre-treatment preferences for Graves. We classify players into two groups based on their pre-disclosure usage of Graves: the first group comprises those who chose Graves in at least $5\%$ of their matches before the coming-out event (henceforth labeled as prior users, $n=953$), while the second group comprises the remaining players (henceforth labeled as non-prior users, $n=6468$). We then examine performance differences both within and between these groups before and after the treatment.
@@
Figure D.II: Players’ average pick rates for Graves and performance measures. Players are divided into two groups based on their preferences for Graves before his disclosure. The panels display the average pick rates for Graves, number of daily matches, and win rate of each group before and after the coming-out event.
@@
The top panel of Figure D.II displays the average pick rate for Graves among prior and non-prior users before and after the treatment. We observe a sharp decline in pick rates among prior users following the coming-out event, similar in size to the decrease shown in Figure 2 and Table 1. Conversely, non-prior users exhibit a marginal increase in average pick rates post-treatment, although this increase is practically negligible.
@@
In the remaining panels of Figure D.II, we investigate whether prior and non-prior users differ in their in-game behavior. First, the bottom left panel shows the average number of daily matches played by each group. We observe similar engagement levels across groups, both before and after the treatment, indicating that prior and non-prior users tend to play a comparable number of matches per day. Notably, neither group appears to disengage from the game following Graves’ disclosure. On the contrary, the number of daily matches slightly increases post-treatment for both groups, suggesting that players are not leaving the game but instead redirecting their attention to other characters. Supporting this interpretation, Figure A.VII in Supplementary Information A shows no meaningful change in role preferences for either group following the event. This pattern is consistent with the idea that players face low switching costs and have access to suitable alternatives, enabling them to express their preferences without disrupting their gameplay. These results align with those displayed in Table 1.
Second, the bottom right panel displays the players’ average win rates. We observe no substantial disparities within or between groups, indicating that differences in pre-disclosure usage of Graves are not associated with differential performance or skill. In particular, the decision to abandon Graves does not coincide with systematic changes in win rates. Overall, these findings provide no evidence that the post-disclosure decline in Graves’ usage is driven by gameplay considerations, thereby strengthening the interpretation that the observed behavior reflects a response to the disclosure rather than changes in in-game incentives.
D.3 Players’ Performance